# LLM-Select: Feature Selection with Large Language Models

### v1

@@ Abstract
In this paper, we demonstrate a surprising capability of large language models (LLMs): given only input feature names and a description of a prediction task, they are capable of selecting the most predictive features, with performance rivaling the standard tools of data science. Remarkably, these models exhibit this capacity across various query mechanisms. For example, we zero-shot prompt an LLM to output a numerical importance score for a feature (e.g., “blood pressure”) in predicting an outcome of interest (e.g., “heart failure”), with no additional context. In particular, we find that the latest models, such as GPT-4, can consistently identify the most predictive features regardless of the query mechanism and across various prompting strategies. We illustrate these findings through extensive experiments on real-world data, where we show that LLM-based feature selection consistently achieves strong performance competitive with data-driven methods such as the LASSO, despite never having looked at the downstream training data. Our findings suggest that LLMs may be useful not only for selecting the best features for training but also for deciding which features to collect in the first place. This could potentially benefit practitioners in domains like healthcare, where collecting high-quality data comes at a high cost.
@@ 1 Introduction
Surprisingly, we find that even without looking at the training data, these models are capable of identifying the most predictive features, with performance rivaling those of data-driven feature selection methods such as the LASSO (Tibshirani, 1996) (see Figure 1(b)). Such a result is counterintuitive; given that datasets with identical input feature names can correspond to vastly different data distributions, due to factors such as selection bias and confounding, it is not obvious that selecting features only based on their names can be effective. Remarkably, we find that for the latest models such as GPT-4, even zero-shot prompting the LLM to generate a numerical feature importance score one-feature-at-a-time can lead to strong feature selection performance. We demonstrate these findings by stress-testing our LLM-based feature selection methods against traditional data-driven methods on a large collection of real-world datasets from various domains, using both closed-source (GPT-4 (OpenAI, 2023), GPT-3.5 (Brown et al., 2020)) and open-source LLMs (Llama-2 (Touvron et al., 2023)) at multiple model scales and applying various prompting strategies.
@@
We propose three approaches to using an LLM for feature selection: (i) selecting features with the highest LLM-generated importance scores, (ii) selecting features based on an LLM-generated ranking, and (iii) sequentially selecting features in a dialogue with an LLM.
@@
We show that LLMs of sufficient scale achieve strong feature selection performance on real-world data, often competitive with traditional data-driven methods such as the LASSO (Tibshirani, 1996).
@@
We comprehensively assess the sensitivity of the LLM-based feature selection methods to various prompting strategies and find that even zero-shot prompting with no additional context about the downstream data can be sufficient for strong feature selection performance.
@@
We show that the feature importance scores generated by larger language models generally exhibit higher correlation with several commonly used feature importance metrics (e.g., Shapley values).
Our findings suggest that LLMs may be useful not only for selecting the most predictive features after data collection, but also for deciding what features to collect in the first place—especially in domains like healthcare and the social sciences where obtaining high-quality data is expensive.
@@ 2.1 Prompting LLMs
Prompting is a computationally efficient and effective method for adapting pretrained LLMs to perform new tasks unseen during training (Radford et al., 2019; Liu et al., 2023). In a standard prompting setup, an output is autoregressively sampled from an LLM, conditional on text descriptions of a desired task and optionally a set of in-context input-output examples (Brown et al., 2020), and used as a solution for the given task. Even without task-specific fine-tuning, such a zero-/few-shot in-context learning approach can be surprisingly effective for adapting pretrained LLMs to solve a wide range of natural-language tasks (Hendrycks et al., 2021; Lin et al., 2022; Patel & Pavlick, 2022; Srivastava et al., 2023), given a language model of sufficient scale (Wei et al., 2022a).
Meanwhile, recent works show that LLM outputs can be highly sensitive to the specifics of the input prompt and the decoding strategy (Jiang et al., 2020; Schick & Schütze, 2021; Zhao et al., 2021) and that choosing an appropriate prompting strategy is crucial, especially for challenging reasoning tasks. In our experiments, we consider two prompting techniques—chain-of-thought prompting (CoT; Wei et al., 2022b) and self-consistency decoding (Wang et al., 2023). We focus on these methods as they often dramatically boost performance on tasks that require multi-step reasoning (Kojima et al., 2022; Lewkowycz et al., 2022; Chen et al., 2023), which we hypothesized to be important for feature selection.
@@
Chain-of-thought prompting (CoT; Wei et al., 2022b) is a few-shot prompting method that augments each input-output example with a chain-of-thought—a coherent series of natural-language reasoning steps that leads to the correct answer. Wang et al. (2023) show that given a large-enough model and an appropriately designed input prompt, CoT prompting can elicit logically consistent step-by-step solutions to unseen examples from the LLM, and substantially boost performance over standard few-shot prompting (Brown et al., 2020) on tasks such as solving math problems (Lewkowycz et al., 2022; Imani et al., 2023), answering commonsense reasoning questions (Suzgun et al., 2023; Anil et al., 2023), and answering expert-level medical questions (Liévin et al., 2022; Singhal et al., 2023)
@@ 2.2 Feature Selection
Feature selection is a classical machine learning problem, where given a set of candidate features, the goal is to select the most informative feature subset that is predictive of an outcome of interest (Blum & Langley, 1997; Guyon & Elisseeff, 2003; Chandrashekar & Sahin, 2014; Li et al., 2017). Feature selection methods can generally be grouped into three categories: filter, wrapper, and embedded methods. Filter methods (Lazar et al., 2012) select features by ranking them according to a statistical or information-theoretic criterion (e.g., mutual information (Lewis, 1992; Ding & Peng, 2005), Fisher score (Duda et al., 2001; Gu et al., 2011), maximum mean discrepancy (Song et al., 2012)) and choosing the top ranked features, independent of the downstream learning algorithm. Wrapper methods (Kohavi & John, 1997) identify a locally optimal feature subset that maximizes the performance of the downstream prediction model using a heuristic search strategy (e.g., sequential selection (Ferri et al., 1994; Luo & Chen, 2014), recursive feature elimination (RFE; Guyon et al., 2002)). Embedded methods select features as part of the model learning process, and are often based on regularization techniques that encourage feature sparsity (Tibshirani, 1996; Yuan & Lin, 2006; Feng & Simon, 2017; Lemhadri et al., 2021). In our experiments, we compare our methods against traditional feature selection baselines from all three categories.
@@
In a prior work most similar to ours, Choi et al. (2022) propose the LMPriors framework, where they prompt the davinci-instruct-beta variant of GPT-3 (Brown et al., 2020) to answer whether each candidate feature should be used to predict the target outcome, and select features whose difference in log-probabilities for generating a “Y” (Yes) or “N” (No) token crosses a predefined threshold. Our work diverges from theirs in two key aspects. First, we propose three different feature selection methods which all directly use the generated text output and not the associated token probabilities, which are often not directly accessible in closed-source, proprietary LLMs. Second, we provide a more comprehensive evaluation across various model scales and prompting strategies on a larger collection of datasets and derive practical insights. In Appendix A.3.3, we also show that our proposed feature selection methods perform as strongly as theirs, even without the same level of access into the LLM.
@@ 3 Selecting Features with LLMs
We address the standard supervised learning setup where, given labeled data with and , our goal is to learn a prediction model such that for some model class and loss function . We assume access to concepts for the input features and for the prediction target, which are text descriptions that capture their real-world semantics. For example, when predicting heart failure (1 if positive, 0 otherwise) given a patient’s blood pressure and weight measurements, , , and denotes the numerical measurements used to learn . Concept annotations are widely available in many practical settings, e.g., as column names in tabular data or via datasheets that contain auxiliary metadata. For feature selection, our goal is to find a subset of size such that a model trained on , where , achieves strong performance under a budget on .
@@ 3.1 Feature Selection with LLMs
To leverage a pretrained LLM for feature selection, we prompt with the input concepts and target concept , and select features based on the generated output. We consider the following three approaches: (i) selecting features based on LLM-generated feature importance scores, (ii) selecting features based on an LLM-generated ranking, and (iii) sequentially selecting features in a dialogue with an LLM. We design separate prompt templates for each approach and denote them by , , and , respectively. Each prompt template can be viewed as a function of the input and target concepts which outputs a set of natural-language instructions tailored to the corresponding selection strategy. While generally, text outputs generated from an LLM need to be processed further to extract the relevant information, we omit such steps in the notation below for simplicity.
@@
We then define to be the set of indices of the top- concepts with the highest importance scores and use for learning . Given that the LLM is only given a single input concept and the target concept each time it is prompted, we hypothesize that captures the marginal importance of each feature for predicting the target, as informed by the knowledge encoded in . We note that the feature importance scores are directly parsed from the text output and do not correspond to the token probabilities associated with generating the text output. We also note that the score range of is an arbitrary choice and therefore evaluate the sensitivity of LLM-Score to different choices in Appendix A.3.4.
@@
We define to be the set of indices of the top- highest ranked concepts and use for learning . We hypothesize that the rank of each input concept reflects its relative importance for predicting the target, with respect to all of the other input concepts in .
@@
In this approach, we consider a selection strategy analogous to sequential selection methods (Ferri et al., 1994; Luo & Chen, 2014). We start with an empty set of concepts and iteratively add a new concept by prompting the LLM to select a candidate concept that would maximally improve cross-validation performance, until we have selected concepts. Formally, at each iteration , we have
where denotes the -th selected input concept, denotes the subset of concept indices selected up to the -th iteration, and . As another setup, we consider starting with containing the concept with the highest score from Equation (1) and iterating over . In the main text, we focus on the former, as it empirically performs better than the latter (we compare the two approaches in Appendix A.3.5). At each , we then use to learn . We hypothesize that this approach encourages selecting a feature that is maximally informative with respect to the feature subset already selected.
@@ 4 Experiments
We include all of the prompt templates that we found to work well in Appendices C–E. In the default template, we only include (i) the main system prompt (e.g., “Your task is to provide a feature importance score between 0 and 1 for predicting target concept and a reasoning behind how the importance score was assigned.”), (ii) output format instructions (e.g., “Output your answer in a JSON format.”), and (iii) the main user prompt (e.g., “Provide a score and reasoning for concept.”). We emphasize that the default prompts are not “fine-tuned” for each dataset in any way, as they only embed the input and target concepts and no other dataset-specific information. Meanwhile, we examine how the following changes to the input prompt affect feature selection performance:
@@
Adding CoT explanations: Given the empirical success of CoT prompting (Wei et al., 2022b) in improving the reasoning capabilities of LLMs in few-shot settings, we investigate whether adding CoT into the few-shot examples (e.g., [“blood pressure”, “Blood pressure is important for… Thus, the score is 0.9.”, 0.9]) improves the performance of LLM-based feature selection.
@@
By default, we use greedy decoding (i.e., sampling with temperature ), given its simple and deterministic behavior222For GPT-4 and GPT-3.5, we average over 5 samples with to account for the inherent non-determinism.. Meanwhile, as there is no clear notion of ground truth when assigning feature importance (e.g., what is the ground-truth importance score for “blood pressure” when predicting “heart failure”?) and multiple reasoning paths are possible for determining importance, we also consider self-consistency decoding (Wang et al., 2023). For the latter, we set and average across 5 samples.
@@ 4.1 Evaluation on Small-Scale Datasets
We compare LLM-Score, LLM-Rank, and LLM-Seq against several feature selection methods using small-scale, low-dimensional datasets from various domains (e.g., healthcare, criminal justice), each with 10–70 features after preprocessing. Here, we focus on the small-scale setting to stress-test the LLM-based feature selection methods in various ways and to allow comparison with methods that are less scalable to high-dimensional settings (e.g., sequential selection baselines). We use seven binary classification datasets (Credit-G, Bank, Give Me Some Credit, COMPAS Recidivism, Pima Indians Diabetes, AUS Cars*, YouTube*) and seven regression datasets (CA Housing, Diabetes Progression, Wine Quality, Miami Housing, Used Cars, NBA*, NYC Rideshare*), where those marked with an asterisk (*) are datasets published after the pretraining-data cutoff dates for all LLMs333For GPT-4 and GPT-3.5, the cutoff date is Sep., 2021 (reference). For Llama-2, the cutoff date is Sep., 2022 (reference).. We include these datasets to address pretraining-data leakage/memorization concerns. We provide full details on all datasets in Appendix A.1.
We evaluate each feature selection method by measuring how the test performance of a downstream prediction model changes as we vary the proportion of features selected from 10% to 100%, in approximately 10% increments. Specifically, for each dataset and at each proportion, we measure the test performance of a downstream -penalized logistic/linear regression model selected via a grid search with 5-fold cross-validation. We use the area under the ROC curve (AUROC) and mean absolute error (MAE) to measure performance on classification and regression tasks, respectively. For LLM-based feature selection, we select the concepts in 10% increments, which may each correspond to more than one feature after preprocessing if the concept is categorical (e.g., one-hot encoding for “ethnicity”).
@@
minimum redundancy maximum relevance selection (MRMR; Ding & Peng, 2005),
@@
For LassoNet and the LASSO, we sweep through a number of regularization strengths to select a variable number of features for downstream training. We then train a separate downstream model as described above, to decouple the effects of feature selection and regularization on test performance. For forward or backward sequential selection, we greedily add/remove a new feature based on the 5-fold cross-validation performance of the downstream prediction model at each iteration. For RFE, we recursively eliminate features with the smallest weights in a logistic/linear regression model selected via a grid search with 5-fold cross-validation using all features. For filtering by mutual information, we select features with the highest marginal mutual information with the target variable, based on the training data. When continuous features and/or labels are present, we use nearest-neighbor approximations (Kraskov et al., 2004; Ross, 2014) to estimate the empirical mutual information. For all experiments, we average the results over 5 random seeds, which control the train-validation splits and random feature selection. We include additional details in Appendix A.2.
@@
Figure 2(a) shows the downstream test performance, averaged over all of the classification (left) and regression (right) datasets, when selecting the top 30% of features according to each baseline and LLM-Score based on GPT-4, GPT-3.5 and Llama-2. For the regression datasets, we report the average ranking based on test MAE to account for differences in the scale of MAE values across datasets. For the larger models, notably GPT-4 (in red) and GPT-3.5 (in green), selecting features based on LLM-generated importance scores leads to strong downstream performance on average, competitive with the data-driven baselines. In Figure A1 of Appendix A.3, we show the same plots for LLM-Rank and LLM-Seq, where we observe similar results. As an example, in Figure 2(b), we show the feature selection paths (i.e., “test performance vs. fraction of features selected” curves) for the best-performing baseline on each dataset, LLM-Score based on GPT-4, and random selection on the YouTube* and NBA* datasets published after the LLM cutoff dates. Here, GPT-4-based LLM-Score outperforms the best baseline (LASSO) on the YouTube* dataset, achieving higher AUROC overall, and performs as strongly as that (forward sequential selection) on the NBA* dataset. Meanwhile, Figure 2(c) shows that performance is less consistent with smaller LLMs. For example, LLM-Score based on Llama-2 (7B) (in purple) performs well on the YouTube* dataset but close to random on the NBA* dataset. These results demonstrate that using LLMs for feature selection can be effective, but suggest that a sufficiently large model may be required for reliable performance. We include the full results on comparing LLM-Score, LLM-Rank, and LLM-Seq to the data-driven baselines in Appendix A.3.1.

### v2

@@ Abstract
In this paper, we demonstrate a surprising capability of large language models (LLMs): given only input feature names and a description of a prediction task, they are capable of selecting the most predictive features, with performance rivaling the standard tools of data science. Remarkably, these models exhibit this capacity across various query mechanisms. For example, we zero-shot prompt an LLM to output a numerical importance score for a feature (e.g., “blood pressure”) in predicting an outcome of interest (e.g., “heart failure”), with no additional context. In particular, we find that the latest models, such as GPT-4, can consistently identify the most predictive features regardless of the query mechanism and across various prompting strategies. We illustrate these findings through extensive experiments on real-world data, where we show that LLM-based feature selection consistently achieves strong performance competitive with data-driven methods such as the LASSO, despite never having looked at the downstream training data. Our findings suggest that LLMs may be useful not only for selecting the best features for training but also for deciding which features to collect in the first place. This could benefit practitioners in domains like healthcare and the social sciences, where collecting high-quality data comes at a high cost.
@@ 1 Introduction
Surprisingly, we find that even without looking at the training data, these models are capable of identifying the most predictive features, with performance rivaling those of data-driven feature selection methods such as the LASSO (Tibshirani, 1996) (Figure 1(b)). Such a result is counterintuitive; given that datasets with identical input feature names can correspond to arbitrarily different data distributions, due to factors such as selection bias and confounding, it is not a priori obvious that selecting features only based on their names will be effective. Remarkably, we find that for the latest models such as GPT-4, even zero-shot prompting the LLM to generate a numerical feature importance score one-feature-at-a-time can lead to strong feature selection performance. Moreover, we find that such LLM-generated feature importance scores are correlated with several commonly used feature importance metrics (e.g., Fisher score (Gu et al., 2011), SHAP (Lundberg & Lee, 2017)), suggesting that LLMs are capable of distilling the real-world relationships encoded in their parameters into statistically meaningful scores. We demonstrate these findings by stress-testing our LLM-based feature selection methods against traditional data-driven methods on real-world datasets from diverse domains (e.g., finance, healthcare, criminal justice), using both closed-source (GPT-4 (OpenAI, 2023), GPT-3.5 (Brown et al., 2020)) and open-source LLMs (Llama-2 (Touvron et al., 2023)) of various sizes.
@@
We propose three approaches to prompting an LLM for feature selection: (i) selecting features with the highest LLM-generated feature importance scores (LLM-Score); (ii) selecting features based on an LLM-generated ranking (LLM-Rank); and (iii) sequentially selecting features in a dialogue with an LLM (LLM-Seq) (Section 3).
@@
We comprehensively assess the sensitivity of LLM-based feature selection methods to various prompting strategies and demonstrate that even zero-shot prompting the LLM with no additional context about the downstream data can elicit strong feature selection performance (Result 3, Section 4.1).
@@
We show that as model scale increases, LLM-generated importance scores generally exhibit higher rank correlation with commonly used feature importance metrics (Result 4, Section 4.1).
Our findings suggest that LLMs may be useful not only for selecting the most predictive features after data collection, but also for deciding what features to collect in the first place. This could benefit practitioners in domains like healthcare and the social sciences, where obtaining high-quality data (e.g., running medical tests for patients, designing survey questions) can be expensive and time-consuming.
@@ 2.1 Prompting LLMs
Prompting is an effective method for adapting off-the-shelf LLMs to perform new tasks unseen during training, without explicit gradient updates to the model parameters (Radford et al., 2019; Liu et al., 2023). In a standard prompting setup, an output is autoregressively sampled from an LLM conditional on text descriptions of a desired task and optionally a set of in-context input-output examples (Brown et al., 2020) and used as a solution for the given task. Even without task-specific fine-tuning, such a zero-shot or few-shot in-context learning approach can be surprisingly effective for adapting pretrained LLMs towards a wide range of natural-language tasks (Hendrycks et al., 2021; Lin et al., 2022; Patel & Pavlick, 2022; Srivastava et al., 2023), given a language model of sufficient scale (Wei et al., 2022a).
Meanwhile, several works show that LLM outputs can be highly sensitive to the specifics of the input prompt and choice of decoding strategy (Jiang et al., 2020; Zhao et al., 2021; Sclar et al., 2024; Jeong et al., 2024a; b). As such, choosing an appropriate prompting strategy is crucial, especially for challenging reasoning tasks. In our experiments, we mainly consider two prompting techniques—chain-of-thought prompting (CoT; Wei et al., 2022b) and self-consistency decoding (Wang et al., 2023). We focus on these methods as they often dramatically boost performance on tasks that require multi-step reasoning (Kojima et al., 2022; Lewkowycz et al., 2022; Chen et al., 2023), which we hypothesized to be important for feature selection.
@@
CoT prompting (Wei et al., 2022b) is a few-shot prompting method that augments each input-output example with a chain-of-thought—a coherent series of natural-language reasoning steps leading to the correct answer. Wang et al. (2023) show that given a large-enough model and an appropriately designed input prompt, CoT prompting can elicit logically consistent step-by-step solutions from the LLM and substantially boost performance on complex tasks such as solving math problems (Lewkowycz et al., 2022; Imani et al., 2023) and answering commonsense (Suzgun et al., 2023; Anil et al., 2023) and knowledge-intensive reasoning questions (Liévin et al., 2022; Singhal et al., 2023).
@@
Extensions of CoT prompting.
More recently, tree-of-thoughts (ToT; Yao et al., 2023) and graph-of-thoughts (GoT; Besta et al., 2024) prompting have been proposed as generalizations of CoT prompting (with self-consistency decoding) to further enhance the exploration and evaluation of multiple reasoning paths. ToT prompting formulates reasoning as a traversal over a tree of plausible partial solutions, allowing for explicit branching and backtracking of diverse reasoning paths in a hierarchical manner. GoT prompting further generalizes ToT prompting by formulating reasoning via a graph, additionally allowing for exploration and dependencies across different reasoning paths. In our paper, we focus on standard CoT prompting (Wei et al., 2022b) for an initial investigation on LLM-based feature selection in the simplest setting. We leave an in-depth investigation of whether extensions such as ToT and GoT prompting can improve LLM-based feature selection performance as future work.
@@ 2.2 Feature Selection
Feature selection is a classical machine learning problem, where given a set of candidate features, the goal is to select the most informative feature subset that is predictive of an outcome of interest (Blum & Langley, 1997; Guyon & Elisseeff, 2003; Chandrashekar & Sahin, 2014; Li et al., 2017). Feature selection methods can generally be grouped into three categories: filter, wrapper, and embedded methods. Filter methods (Lazar et al., 2012) select features by ranking them according to a statistical or information-theoretic criterion—e.g., mutual information (Lewis, 1992; Ding & Peng, 2005; Bennasar et al., 2015), Fisher score (Duda et al., 2001; Gu et al., 2011), maximum mean discrepancy (Song et al., 2012)—and choosing the top ranked features, independent of the downstream learning algorithm. Wrapper methods identify a locally optimal feature subset that maximizes the performance of the downstream prediction model (Kohavi & John, 1997; Yamada et al., 2014; Chen et al., 2017; Climente-González et al., 2019), often by employing a heuristic search strategy (e.g., sequential selection (Ferri et al., 1994; Luo & Chen, 2014; Yasuda et al., 2023), recursive feature elimination (RFE; Guyon et al., 2002)). Embedded methods select features as part of the model learning process, most commonly based on regularization techniques that encourage feature sparsity (Tibshirani, 1996; Yuan & Lin, 2006; Feng & Simon, 2017; Lemhadri et al., 2021), and others based on specialized neural network architectures (Lu et al., 2018; Balın et al., 2019; Yamada et al., 2020). In our experiments, we compare our methods against traditional feature selection baselines from all three categories.
@@
In a prior work most similar to ours, Choi et al. (2022) propose the LMPriors framework, where they prompt the davinci-instruct-beta variant of GPT-3 (Brown et al., 2020) to answer whether each candidate feature should be used to predict the target outcome, and select features whose difference in log-probabilities for generating a “Y” (Yes) or “N” (No) token crosses a predefined threshold. Our work differs from theirs in two key aspects. First, we propose three different feature selection methods which all directly use the generated text output and not the associated token probabilities, which are often not directly accessible in closed-source, proprietary LLMs. Second, we provide a more comprehensive evaluation across various model scales and prompting strategies on a larger collection of datasets and derive practical insights. Our proposed feature selection methods perform as strongly as theirs even without the same level of access into the LLM (Appendix A.3.3).
@@ 3 Selecting Features with LLMs
We address the standard supervised learning setup where, given labeled data with and , our goal is to learn a prediction model such that for some model class and loss function . We assume access to concepts for the input features and for the prediction target, which are text descriptions that capture their real-world semantics. For example, when predicting heart failure (1 if positive, 0 otherwise) given a patient’s blood pressure and weight measurements, , , and denotes the numerical measurements used to learn . Concept annotations are widely available in many practical settings, e.g., as column names in tabular datasets or via datasheets that contain auxiliary metadata. For feature selection, our goal is to find a subset of size such that a model trained on , where , achieves strong performance under a budget on .
@@ 3.1 Feature Selection with LLMs
To leverage a pretrained LLM for feature selection, we prompt with the input concepts and target concept , and select features based on the generated output. We consider the following three approaches: (i) selecting features based on LLM-generated feature importance scores; (ii) selecting features based on an LLM-generated ranking; and (iii) sequentially selecting features in a dialogue with an LLM. We design separate prompt templates for each approach and denote them by , , and , respectively. Each prompt template can be viewed as a function of the input and target concepts which outputs a set of natural-language instructions tailored to the corresponding selection strategy. While generally, text outputs generated from an LLM need to be processed further to extract the relevant information, we omit such steps in the notation below for simplicity.
@@
We then define to be the set of indices of the top- concepts with the highest importance scores and use to learn a downstream prediction model . Given that the LLM is only given a single input concept and the target concept each time it is prompted, we hypothesize that captures the marginal importance of each feature for predicting the target, as informed by the knowledge encoded in . We note that the feature importance scores are directly parsed from the text output and do not correspond to the token probabilities associated with generating the text output. We also note that the score range of is an arbitrary choice and therefore evaluate the sensitivity of LLM-Score to different choices in Appendix A.3.4.
@@
We define to be the set of indices of the top- highest ranked concepts and use to learn a downstream prediction model . We hypothesize that the rank of each input concept reflects its relative importance for predicting the target, with respect to all of the other input concepts in .
@@
In this approach, we consider a selection strategy analogous to sequential selection methods. We start with an empty set of concepts and iteratively add a new concept by prompting the LLM to select a candidate concept that would maximally improve the cross-validation performance of a downstream prediction model. Formally, assuming that our goal is to select concepts, at each iteration , we have
where denotes the -th selected input concept, denotes the subset of concept indices selected up to the -th iteration, and . We then use to train a downstream prediction model , where we tune the hyperparameters via 5-fold cross-validation. The cross-validation performance of on is then appended to the prompt used for the next iteration (Appendix E). For the initialization of the feature subset, we also consider starting with containing the concept with the highest score from Equation (1) and iterating over . However, we focus on the former approach in the main text, as it empirically performs better than the latter (we compare the two approaches in Appendix A.3.5). We hypothesize that this approach encourages the LLM to select a feature that is maximally informative with respect to the feature subset already selected, at each iteration of the algorithm. Meanwhile, we note that since LLM-Seq is a greedy sequential selection approach, the resulting top- feature subset is not guaranteed to be globally optimal. This limitation arises as at each iteration , the next feature to add is determined with respect to , which precludes an exploration all possible subsets of .
@@ 4 Experiments
We provide all of the prompt templates used in our experiments in Appendices C–E. The prompt templates were carefully constructed to reliably elicit the desired response in the correct format from each LLM. In the default template, we only include (i) the main system prompt (e.g., “Your task is to provide a feature importance score between 0 and 1 for predicting target concept and a reasoning behind how the importance score was assigned.”), (ii) output format instructions (e.g., “Output your answer in a JSON format.”), and (iii) the main user prompt (e.g., “Provide a score and reasoning for concept.”). We emphasize that the default prompts are not “fine-tuned” on each dataset in any way, as they only embed the input and target concepts and no other dataset-specific information. Meanwhile, we examine how the following changes to the input prompt affect feature selection performance:
@@
Adding CoT explanations: Given the empirical success of CoT prompting (Wei et al., 2022b) in improving the reasoning capabilities of LLMs in few-shot settings, we investigate whether adding CoT reasoning into the few-shot examples (e.g., [“blood pressure”, “Blood pressure is important for Thus, the score is 0.9.”, 0.9]) improves the performance of LLM-based feature selection.
For all three changes, we manually construct the relevant inputs via human annotation. For dataset-specific context, we manually summarize the metadata associated with each dataset (e.g., sourced from the “Data Card” of a Kaggle dataset). For selecting the few-shot examples and generating their CoT explanations, we prioritize features whose semantic relevance to the target is unambiguous and least open to subjective interpretation, based on common sense and relevant domain knowledge. For a given dataset, the few-shot examples and CoT explanations provided to the LLMs remain fixed across all queries. See Appendix C for full details on the examples and CoT explanations for each dataset considered in Sections 4.1–4.2.
@@
By default, we use greedy decoding (i.e., sampling with temperature ), given its straightforward and deterministic behavior. For GPT-4 and GPT-3.5, we account for their inherent non-determinism at the time of writing by sampling 5 responses with and averaging the feature importance scores parsed from the text generations. Meanwhile, as there is no clear notion of ground truth when assigning feature importance (e.g., what is the ground-truth importance score for “blood pressure” when predicting “heart failure”?) and multiple reasoning paths are possible for determining importance, we also consider self-consistency decoding (Wang et al., 2023). For the latter, we set and average across 5 samples.
Source code.
To ensure the reproducibility of our results, we open-source the source code used for all of our evaluations detailed below via our GitHub repository222https://github.com/taekb/llm-select.
@@ 4.1 Evaluation on Small-Scale Datasets
We compare LLM-Score, LLM-Rank, and LLM-Seq against several feature selection methods using small-scale, low-dimensional datasets from various domains (e.g., healthcare, criminal justice), each with 10–70 features after preprocessing. Here, we focus on the small-scale setting to stress-test the LLM-based feature selection methods in various ways and to allow comparison with methods that are less scalable to high-dimensional settings (e.g., sequential selection baselines). We use seven binary classification datasets (Credit-G, Bank, Give Me Some Credit, COMPAS Recidivism, Pima Indians Diabetes, AUS Cars*, YouTube*) and seven regression datasets (CA Housing, Diabetes Progression, Wine Quality, Miami Housing, Used Cars, NBA*, NYC Rideshare*), where those marked with an asterisk (*) are datasets published after the pretraining data cutoff dates of all of the LLMs we evaluate333For GPT-4 and GPT-3.5, the cutoff date is Sep., 2021 at the time of writing (reference). For Llama-2, the cutoff date is Sep., 2022 (reference). As LLMs are regularly updated, the knowledge cutoff dates may be different in more recent versions.. We include these datasets to ensure that our findings generalize to datasets that the LLMs have not been trained on. We provide the remaining details on all datasets in Appendix A.1.
We evaluate each feature selection method by measuring how the test performance of a downstream prediction model changes as we vary the proportion of features selected from 10% to 100%, in approximately 10% increments. On each dataset and at each proportion, we measure the test performance of an -penalized logistic/linear regression model trained using the selected features. For each training run, we perform model selection via grid search with 5-fold cross-validation. We use the area under the ROC curve (AUROC) and mean absolute error (MAE) to measure performance on classification and regression tasks, respectively. For LLM-based feature selection, we select the concepts in 10% increments, which may each correspond to more than one feature after preprocessing if the concept is categorical (e.g., one-hot encoding for “ethnicity”).
@@
minimum redundancy maximum relevance (MRMR; Ding & Peng, 2005),
@@
HSIC-Lasso (Yamada et al., 2014),
9.
Concrete Autoencoder (CAE; Balın et al., 2019),
10.
Sequential Attention (SA; Yasuda et al., 2023),
11.
@@
For LassoNet and the LASSO, we first compute the regularization paths with warm starts (Friedman et al., 2010) to identify the regularization coefficients corresponding to each feature proportion, and then train a separate downstream model as described above. We take this two-step approach to decouple the effects of feature selection and regularization on the downstream test performance. For forward/backward sequential selection, we greedily add/remove a new feature at each iteration based on the 5-fold cross-validation performance resulting from adding/removing each candidate feature. For RFE, we recursively eliminate features with the smallest weights in a logistic/linear regression model selected via a grid search with 5-fold cross-validation using all features. For MI, we select features with the highest marginal mutual information with the target variable based on the training data. When continuous features and/or labels are present, we use the nearest-neighbor approximations (Kraskov et al., 2004; Ross, 2014) available in scikit-learn to estimate the empirical mutual information. For HSIC-Lasso, we compute the full kernel matrix if the number of training examples is less than 1000, but otherwise use the block-wise approximation method by Climente-González et al. (2019), using the recommended block size and number of permutations . For CAE, we use a multi-layer perceptron (MLP) decoder444We use the architecture in the official repository for CAE: https://github.com/mfbalin/Concrete-Autoencoders. with hidden layers of width 256 and train the model for a maximum of 1000 epochs with a supervised learning objective (as in Appendix F of Balın et al. (2019)). For SA, we use an MLP with 1 hidden layer of width 67 and ReLU activation, following Yasuda et al. (2023). For all experiments, we repeat and average the results over 5 random seeds, which control the train-validation splits used for cross-validation and the behavior of random feature selection, to ensure the robustness of results. We include the remaining details in Appendix A.2.
@@
Figure 2(a) shows the downstream test performance, averaged over all of the classification (left) and regression (right) datasets, when selecting the top 30% of features according to each baseline and LLM-Score based on GPT-4, GPT-3.5 and Llama-2. For the regression datasets, we report the average ranking based on test MAE to account for differences in the scale of MAE values across datasets. For the larger models, notably GPT-4 (in blue) and GPT-3.5 (in green), selecting features based on LLM-generated importance scores leads to strong downstream performance on average, competitive with the data-driven baselines. We observe similar results for LLM-Rank and LLM-Seq (Figure A1). As an example, in Figure 2(b), we show the feature selection paths (i.e., “test performance vs. fraction of features selected” curves) for the best-performing baseline on each dataset, LLM-Score based on GPT-4, and random selection on the YouTube* and NBA* datasets published after the LLM cutoff dates. Here, GPT-4-based LLM-Score outperforms the best baseline (LASSO) on the YouTube* dataset, achieving higher AUROC overall, and performs as strongly as that (HSIC-Lasso) on the NBA* dataset. Meanwhile, performance is less consistent with smaller LLMs (Figure 2(c)). For example, LLM-Score based on Llama-2 (7B) (in yellow) performs well on the YouTube* dataset but close to random on the NBA* dataset. These results demonstrate that using LLMs for feature selection can be effective, but suggest that a sufficiently large model may be required for reliable performance. We include the full results on comparing LLM-Score, LLM-Rank, and LLM-Seq to the data-driven baselines in Appendix A.3.1.